# Polar Sparsity: High Throughput Batched LLM Inferencing with Scalable Contextual Sparsity

### v2

@@ 5.1 Benchmark Evaluation
Figure 4 shows the zero-shot accuracy on downstream tasks as we vary the attention density. At each data point, we activate the attention heads/groups with the highest output logits as predicted by the routers. For OPT models, we also apply dynamic layer-wise sparsity to MLP layers. Across all models, we observe that most tasks can be solved accurately even under high attention head sparsity with minimal degradation up to a critical threshold. The evaluation results show that this critical point varies with the architecture and size of the model. Stable average accuracy is maintained down to 30% attention density for OPT-66b, 50% for smaller OPT-6.7b, LLaMA-2 7b/13b models, and 62.5% for GQA models like LLaMA 3.1 70b. GQA models have naturally less KV diversity in the attention layers and group sparsity is also inherently weaker compared to head sparsity, as it may overlook important heads that reside in inactive groups. This limitation likely contributes to the faster accuracy degradation observed in LLaMA 3.1 70b model.
Furthermore, we observe that some challenging tasks like RTE and MMLU are more sensitive to head sparsity and require more active heads. This task-dependent sensitivity is particularly encouraging, as it suggests that most tasks can be served with only the most critical heads, while harder tasks could potentially activate more for higher accuracy within the same batch. Given that head sparsity is batch-invariant, this opens the door to context-aware, per-sequence head activation, paving the way for lossless sparse inference, a direction we discuss in future work. Table 1 presents the accuracy of the sparsified models at their respective critical attention density thresholds.
@@ 5.2 Generation Throughput
Figure 5: OPT models sparse decoding throughput, seq len 1920 using pipeline parallelism. (a) OPT 6.7b, critical threshold 50%. (b) OPT 66b, critical threshold 30%. Polar Sparsity delivers up to $\times2$ higher throughput than dense and up to 2×2$\times$2 × than standard activation sparsity at scale.
@@
Figure 6: LLaMA models sparse decode throughput results using pipeline parallelism. (a) LLaMA-2-7b seq len 3968, critical threshold 50%.(b) LLaMA-3.1-70b seq len 8192, critical threshold 62.5%. Polar Sparsity delivers up to 1.85×1.85$\times$1.85 × higher throughput than dense baseline at scale.
@@
Figure 5 shows the throughput results of OPT models using Deja Vu-style Dejavu activation sparsity and Polar Sparsity with the dense baseline at different batch sizes. We substitute DejaVu’s sparse GEMV with our Selective GEMM with dynamic top-$k$ to support efficient batched execution. At batch size 1, Polar Sparsity performance is similar to that of conventional activation sparsity. As batch size increases, the effectiveness of conventional activation sparsity diminishes due to the reduced union sparsity across sequences. In contrast, Polar Sparsity leverages the stable, batch-invariant sparsity of attention heads and scales efficiently. Polar Sparsity delivers up to $\times2$ higher decoding throughput than dense and 2×2$\times$2 × more than DejaVu at large batch sizes. In LLaMA-2 models, where sparsity is applied only to attention layers, speedups become apparent at larger batch sizes, reaching up to 1.85×1.85$\times$1.85 ×. Figure 6 presents throughput results for LLaMA 3.1 70B. Despite higher attention density (62.5%), Polar Sparsity achieves a 1.51×1.51$\times$1.51 × speedup at large batch sizes. While we primarily report throughput, Polar Sparsity naturally reduces inter-token latency by a similar margin and accelerates per-query generation. Appendix E provides a detailed evaluation across models and workloads.
@@ 6 Limitations and Future Work
Polar Sparsity is most effective in large-scale, high-throughput settings, particularly during batched decoding. Its benefits diminish for smaller models or small-batch inference due to limited GPU workload and reduced parallelism. Additionally, we fix the top-k activated heads/groups per layer, whereas a dynamic, input- or layer-adaptive strategy could yield better efficiency and accuracy. Head sparsity could also be combined with token sparsity for potential multiplicative gains. We expect our approach to provide efficiency gains in attention variants like Multi-Query mqa and Multi-Latent Attention deepseekv2 ; however, the critical threshold could be higher similar to GQA.
Our approach maintains average accuracy within 1% of the original model at the critical threshold, and this minor drop could potentially be recovered through targeted fine-tuning. Further gains may be possible by training models to induce higher attention head or group sparsity. Since head/group sparsity is batch invariant, there is a promising opportunity to explore task-aware activation, allocating more heads to harder queries and fewer to easier ones within the same batch. Moreover, decoding difficulty can vary across tokens, even within a single sequence. Some tokens may be predicted with fewer active heads, allowing higher sparsity to be applied adaptively at each decoding step. A fine-grained router that dynamically selects heads based on context and difficulty could unlock lossless sparse inference with even greater throughput. We see this as a promising direction toward more adaptive, efficient, and task-aware LLM inference systems.
@@ 7 Conclusion
Our work highlights the scalability and effectiveness of contextual sparsity for accelerating batched LLM inference. We introduce Polar Sparsity, a key insight showing that as batch size and sequence length grow, the importance of sparsity transitions from MLP layers, where union activation increases, to Attention layers, where head-level sparsity remains stable and batch-invariant. By developing sparsity-aware GPU kernels for both MLP and Attention layers, we achieve consistent speedups across a wide range of models, batch sizes and sequence lengths with minimal impact on accuracy. Our results are competitive with state-of-the-art approaches and delivers up to $\times2$ end-to-end speedups in large-scale settings, affirming the practical viability of Polar Sparsity for efficient and scalable LLM serving. This method is a step towards realizing scalable, high-performance, batched LLM inference that meets the growing demands of modern applications.
@@ Acknowledgments
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 (LLNL-CONF-2005579). This research was also supported in part by the National Science Foundation under Grant No. 2203033.
@@ Appendix A Extended Related Works
Inference efficiency has been a central focus in machine learning systems, with a diverse set of strategies such as sparsity, pruning, quantization, distillation, speculative decoding, SSD-based offloading, and advanced memory management techniques proposed to reduce computational cost, memory footprint, and latency [7, 35, 25, 31, 23, 48]. These techniques address distinct bottlenecks and are often complementary, enabling combined use for greater gains.
@@ A.1 Contextual Activation Sparsity
Recent work on activation sparsity in large language models has explored a variety of mechanisms to reduce inference costs without sacrificing accuracy. Early studies revealed that only a small fraction of neurons in transformer MLP layers are active for any given input, establishing the foundation for sparsity-aware inference [33]. Building on this, ReLUfication and ProSparse showed that replacing GELU or SwiGLU with ReLU or progressively encouraging ReLU-like behavior exposes inherent sparsity and improves inference efficiency, even without retraining [41, 50].
Training-free methods such as TEAL and R-Sparse apply post-hoc magnitude-based pruning across models like LLaMA-2 and LLaMA-3 and preserve accuracy [34, 61]. CATS improves on this by making the threshold contextually adaptive, dynamically pruning low-activation neurons based on input features [30]. Deja Vu and ShadowLLM train lightweight predictors to anticipate important neurons and attention heads token-by-token, enabling dynamic sparsity at inference time [36, 3]. Our work draws inspiration from previous predictive approaches. Complementary to predictor-based sparsity, LTE and GRIFFIN optimize structured sparsity patterns during pretraining or in a prompt-aware manner, enabling the model to internalize which parts of the architecture to use for different inputs [62, 15]. Finally, hardware-aware methods like PowerInfer leverage the heavy-tailed distribution of neuron activations to map “hot” neurons onto GPU memory while streaming “cold” ones from the CPU, delivering efficient offloading based inference [51].
While prior work demonstrates promising gains in activation sparsity, most methods primarily target single-query inference and small batch sizes, where neuron sparsity remains high. However, these approaches often fail to scale under batching, as the union of active neurons across inputs approaches dense inference, diminishing the benefits of sparsity. In this work, we build on these ideas and extend activation sparsity to the large-batch regime by leveraging the batch-invariant sparsity patterns of attention heads. We design hardware-efficient GPU kernels that exploit this structure, enabling scalable and high-throughput LLM inference with significant efficiency gains at larger batch sizes.
@@ A.2 Attention Head Sparsity and Token Sparsity
Several recent works explore token sparsity to reduce KV cache and attention computation during generative inference. HashAttention introduces semantic sparsity by mapping tokens to a learned Hamming space, enabling efficient selection of pivotal tokens using hash collisions [13]. A2SF proposes accumulative attention scoring with a forgetting factor to fairly rank token importance across time steps in autoregressive decoders [44]. Keyformer and VATP both address token pruning by extending beyond attention scores: Keyformer selects a small set of “key tokens” based on attention concentration [1] , while VATP incorporates the value vector norm into token importance estimation [20]. Native Sparse Attention introduces a hardware-aligned sparse attention mechanism with trainable sparsity patterns suited for long-sequence modeling [57], and MoBA applies mixture-of-expert routing principles to attention tokens by dynamically selecting relevant blocks for each query [37]. These approaches are largely orthogonal to our method and can potentially be combined with Polar Sparsity for multiplicative gains.
Two recent methods, Mixture of Attention Heads (MoA) [59] and MoH: Mixture-of-Head Attention [26], treat attention heads as experts and learn routers to activate a subset per token. MoA requires full model training from scratch, while MoH involves extensive fine-tuning of pre-trained models to learn dynamic head weighting. Although both demonstrate accuracy improvements, they offer limited efficiency gains due to reliance on dense attention kernels. In contrast, our method keeps the model backbone fixed, trains only lightweight routers for head activation, and leverages our custom Select Head Attention kernel to achieve real wall-clock speedups. We believe prior work was limited by the absence of efficient sparse attention primitives, and with the integration of our kernel, these and future approaches have the potential to achieve both high accuracy and meaningful efficiency gains. We hope this unlocks new opportunities and renews momentum in head sparsity research for accelerating LLM inference.
@@ B.1 MLP Activations
(b) OPT 30b
Figure 7: OPT Family Neuron Batch Activations
(a) ReLU Llama-2-7b
@@
Figure 8: Llama-2-7b Sparsified Models Neuron Batch Activations
@@
OPT models, which use ReLU activations, exhibit strong inherent sparsity. As shown in Figure 7, early layers are significantly sparser than deeper ones, a trend consistent across all OPT models. While increasing batch size reduces overall sparsity, the early layers retain enough sparsity to benefit from selective execution. To test the limits of sparsity, we simulated a batch of 10,000 samples and observed that the activations in the initial layers do approach dense and are not, in fact, dead neurons. However, for practical batch sizes, these layers are sufficiently sparse to support accelerated inference. Recent work has shown that sparsity in ReLU-based models tends to increase during pretraining [38]. The OPT models, having undergone extensive large-scale pretraining, likely benefit from this effect, resulting in more structured and persistent activation sparsity.
In contrast, sparsity patterns in sparsified LLaMA-2-7B models are more sensitive to batching. Figure 8 shows that replacing SwiGLU with ReLU yields high initial sparsity that quickly diminishes with batching and lacks a clear layerwise trend. Figure 8 shows that Prosparse introduces deeper-layer sparsity, but with high variance and similarly degrading trends at larger batch sizes. Unlike OPT, these Sparse-LLaMA variants are only lightly fine-tuned on smaller datasets, which likely limits the emergence of stable sparsity patterns observed in fully pretrained models. Given these observations, we focus our selective execution in LLaMA models on attention head sparsity, where the activation structure remains more consistent across batches.
@@ Appendix C Sparsity Prediction
Our MLP router follows the design used in prior contextual sparsity work [36, 51]. Each router is a two-layer feedforward network with a hidden dimension of 1024, trained independently for each transformer layer. To partially hide the latency of the MLP router, we overlap its execution with the Attention layer. It is trained using supervised learning with ground-truth neuron activations derived from dense MLP outputs. At inference time, the router produces scalar logits for each neuron, which are used to rank and select the top-$k$ neurons for activation. The attention head routers use a single layer feedforward network as described in Section 4.2. Since the attention routers are much smaller than the MLP routers, we simply run them synchronously before each attention layer. The routers are optimized as binary classifiers using a binary cross-entropy loss, with the AdamW optimizer. We use a batch size of 64, a learning rate of 1e-4, and early stopping over a maximum of 20 epochs. The LLM parameters are frozen during router training. Supervision data is collected from inference runs on the WikiText-2 dataset, as described in Section 5. To determine minimal top-k values for the MLP layers, we apply a simple greedy algorithm (Algorithm 2) that incrementally adjusts the threshold to meet the target recall of 99%. This calibration is performed offline for each OPT model variant.
@@
5:     $\hat{A}\leftarrow R(H)$triangleright Predict activations

### v3

@@ 5.1 Benchmark Evaluation
Figure 5 shows the zero-shot accuracy on downstream tasks as we vary the attention density. At each data point, we activate the attention heads/groups with the highest output logits as predicted by the routers. For OPT models, we also apply dynamic layer-wise sparsity to MLP layers. Across all models, we observe that most tasks can be solved accurately even under high attention head sparsity with minimal degradation up to a critical threshold. The evaluation results show that this critical point varies with the architecture and size of the model. Stable average accuracy is maintained down to 30% attention density for OPT-66b, 50% for smaller OPT-6.7b, LLaMA-2 7b/13b models, and 62.5% for GQA models like LLaMA 3.1 70b.
Table 1 and 2 presents the accuracy of the sparsified models at their respective critical attention density thresholds. Polar Sparsity achieves performance comparable to the dense baseline, with average accuracy differences within 1%. Instruction-tuned models also sustain strong performance on LongBench under Polar Sparsity, underscoring its effectiveness on challenging generative benchmarks.
We observe that some challenging tasks like RTE, MMLU, MMLU PRO are more sensitive to head sparsity and require more active heads. This task-dependent sensitivity is particularly encouraging, as it suggests that most tasks can be served with only the most critical heads, while harder tasks could potentially activate more for higher accuracy within the same batch. Given that head sparsity is batch-invariant, this opens the door to context-aware, per-sequence head activation, paving the way for lossless sparse inference, a direction we discuss in future work.
@@
Table 2: Evaluation of instruction-tuned LLMs at critical sparsity thresholds. Polar Sparsity maintains strong performance on generative tasks.
Model MMLU PRO LongBench-e
Mistral-7B-inst 0.247 0.392
Mistral-7B-inst + PolarSparse-0.5 0.244 0.388
Llama-3.1-8B-inst 0.409 0.443
Llama-3.1-8B-inst + PolarSparse-0.625 0.384 0.429
Qwen-2.5-14B-inst 0.497 0.421
Qwen-2.5-14B-inst + PolarSparse-0.625 0.457 0.414
@@ 5.2 Generation Throughput
Figure 6: OPT models sparse decoding throughput, seq len 1920 using pipeline parallelism. (a) OPT 6.7b, critical threshold 50%. (b) OPT 66b, critical threshold 30%. Polar Sparsity delivers up to $2.2\times$ higher throughput than dense and up to $2\times$ than standard activation sparsity at scale.
@@
Figure 7: LLaMA models sparse decode throughput results using pipeline parallelism. (a) LLaMA-2-7b seq len 3968, critical threshold 50%.(b) LLaMA-3.1-70b seq len 8192, critical threshold 62.5%. Polar Sparsity delivers up to $1.85\times$ higher throughput than dense baseline at scale.
@@
Figure 6 shows the throughput results of OPT models using Deja Vu-style Dejavu activation sparsity and Polar Sparsity with the dense baseline at different batch sizes. We substitute DejaVu’s sparse GEMV with our Selective GEMM with dynamic top-$k$ to support efficient batched execution. At batch size 1, Polar Sparsity performance is similar to that of conventional activation sparsity. As batch size increases, the effectiveness of conventional activation sparsity diminishes due to the reduced union sparsity across sequences. In contrast, Polar Sparsity leverages the stable, batch-invariant sparsity of attention heads and scales efficiently. Polar Sparsity delivers up to $2.2\times$ higher decoding throughput than dense and $2\times$ more than DejaVu at large batch sizes. In LLaMA-2 models, where sparsity is applied only to attention layers, speedups become apparent at larger batch sizes, reaching up to $1.85\times$. Figure 7 presents throughput results for LLaMA 3.1 70B. Despite higher attention density (62.5%), Polar Sparsity achieves a $1.51\times$ speedup at large batch sizes. While we primarily report throughput, Polar Sparsity naturally reduces inter-token latency by a similar margin and accelerates per-query generation. Appendix E provides a detailed evaluation across models and workloads.
@@ 6 Limitations and Future Work
Polar Sparsity is most effective in large-scale, high-throughput settings, particularly during batched decoding. Its benefits diminish for small-batch inference due to limited GPU workload and reduced parallelism. Additionally, we fix the top-k activated heads/groups per layer, whereas a dynamic, input- or layer-adaptive strategy could yield better efficiency and accuracy. Head sparsity could also be combined with token sparsity for potential multiplicative gains. Our experiments use benchmarks with context lengths up to 16k tokens. As newer LLMs reach million-token contexts, extending our analysis to these ultra-long settings is an important direction for future work. We expect our approach to provide efficiency gains in attention variants like Multi-Query mqa and Multi-Latent Attention deepseekv2; however, the critical threshold could be higher similar to GQA. GQA models have naturally less KV diversity in the attention layers and group sparsity is also inherently weaker compared to head sparsity, as it may overlook important heads that reside in inactive groups. This limitation likely contributes to the faster accuracy degradation observed in GQA models. Our current evaluation focuses on greedy decoding. Extending Polar Sparsity to beam search and speculative decoding, where sparsity patterns may differ, offers an exciting direction for future research.
Our approach maintains average accuracy within 1% of the original model at the critical threshold, and this minor drop could potentially be recovered through targeted fine-tuning. Further gains may be possible by training models to induce higher attention head or group sparsity. Since head/group sparsity is batch invariant, there is a promising opportunity to explore task-aware query-sensitive routing, allocating more heads to harder queries and fewer to easier ones within the same batch. Moreover, decoding difficulty can vary across tokens, even within a single sequence. Some tokens may be predicted with fewer active heads, allowing higher sparsity to be applied adaptively at each decoding step. A fine-grained router that dynamically selects heads based on context and difficulty could unlock lossless sparse inference with even greater throughput. We see this as a promising direction toward more adaptive, efficient, and task-aware LLM inference systems.
@@ 7 Conclusion
Our work highlights the scalability and effectiveness of contextual sparsity for accelerating batched LLM inference. We introduce Polar Sparsity, a key insight showing that as batch size and sequence length grow, the importance of sparsity transitions from MLP layers, where union activation increases, to Attention layers, where head-level sparsity remains stable and batch-invariant.To exploit this property, we develop Selective Head Attention with sparsity-aware GPU kernels that execute computations only for activated heads and neurons. Together, these optimizations deliver consistent speedups across a wide range of models, batch sizes, and sequence lengths with minimal impact on accuracy. Our results are competitive with state-of-the-art approaches and delivers up to $2.2\times$ end-to-end speedups in large-scale settings, affirming the practical viability of Polar Sparsity for efficient and scalable LLM serving. This method is a step towards realizing scalable, high-performance, batched LLM inference that meets the growing demands of modern applications.
@@ Acknowledgments
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 (LLNL-CONF-2005579). This research was also supported in part by the National Science Foundation under Grant No. 2203033. A portion of this work was conducted during an internship at NVIDIA.
@@ Appendix A Extended Related Works
Inference efficiency has been a central focus in machine learning systems, with a diverse set of strategies such as sparsity, pruning, quantization, distillation, speculative decoding, SSD-based offloading, and advanced memory management techniques proposed to reduce computational cost, memory footprint, and latency [efficientQAT, vptq, distillingstepbystep, speculativedecoding, gammatune, knowledgedistillation, espn+, vLLM]. These techniques address distinct bottlenecks and are often complementary, enabling combined use for greater gains.
@@ A.1 Contextual Activation Sparsity
Recent work on activation sparsity in large language models has explored a variety of mechanisms to reduce inference costs without sacrificing accuracy. Early studies revealed that only a small fraction of neurons in transformer MLP layers are active for any given input, establishing the foundation for sparsity-aware inference [lazyneuron]. Building on this, ReLUfication and ProSparse showed that replacing GELU or SwiGLU with ReLU or progressively encouraging ReLU-like behavior exposes inherent sparsity and improves inference efficiency, even without retraining [relufication, prosparse].
Training-free methods such as TEAL and R-Sparse apply post-hoc magnitude-based pruning across models like LLaMA-2 and LLaMA-3 and preserve accuracy [tealsparsity, rsparse]. CATS improves on this by making the threshold contextually adaptive, dynamically pruning low-activation neurons based on input features [catsthreshold]. Deja Vu and ShadowLLM train lightweight predictors to anticipate important neurons and attention heads token-by-token, enabling dynamic sparsity at inference time [Dejavu, shadowllm]. Our work draws inspiration from previous predictive approaches. Complementary to predictor-based sparsity, LTE and GRIFFIN optimize structured sparsity patterns during pretraining or in a prompt-aware manner, enabling the model to internalize which parts of the architecture to use for different inputs [learnefficient, griffin]. Finally, hardware-aware methods like PowerInfer leverage the heavy-tailed distribution of neuron activations to map “hot” neurons onto GPU memory while streaming “cold” ones from the CPU, delivering efficient offloading based inference [powerinfer].
Existing methods for improving batch efficiency using contextual sparsity are limited in scope. PowerInfer considers batch sizes up to 32 but targets CPU-based offloading systems, where the baseline is already constrained by CPU and PCIe bandwidth [powerinfer]. Herd attempts to group sequences with similar sparsity patterns to increase overlap [herd], but operates at batch sizes $\leq$4 and is difficult to apply in dynamic, real‑world random queries, where identifying on‑the‑fly sequences that share similar activation patterns becomes prohibitively challenging. These approaches sparsify only the MLP block and speedup decreases with larger batch sizes. In contrast, our method handles arbitrary sequences and scales effectively to large batch sizes, aligning with realistic deployment scenarios.
@@ A.2 Attention Head Sparsity and Token Sparsity
Several recent works explore token sparsity to reduce KV cache and attention computation during generative inference. HashAttention introduces semantic sparsity by mapping tokens to a learned Hamming space, enabling efficient selection of pivotal tokens using hash collisions [hashattention]. A2SF proposes accumulative attention scoring with a forgetting factor to fairly rank token importance across time steps in autoregressive decoders [a2sfaccumulativeattentionscoring]. Keyformer and VATP both address token pruning by extending beyond attention scores: Keyformer selects a small set of “key tokens” based on attention concentration [keyformer] , while VATP incorporates the value vector norm into token importance estimation [attentionscoreisnotallyouneed]. Native Sparse Attention introduces a hardware-aligned sparse attention mechanism with trainable sparsity patterns suited for long-sequence modeling [NSA], and MoBA applies mixture-of-expert routing principles to attention tokens by dynamically selecting relevant blocks for each query [MoBA]. These approaches are largely orthogonal to our method and can potentially be combined with Polar Sparsity for multiplicative gains.
Two recent methods, Mixture of Attention Heads (MoA) [MoA] and MoH: Mixture-of-Head Attention [MoH], treat attention heads as experts and learn routers to activate a subset per token. MoA requires full model training from scratch, while MoH involves extensive fine-tuning of pre-trained models to learn dynamic head weighting. Although both demonstrate accuracy improvements, they offer limited efficiency gains due to reliance on dense attention kernels. In contrast, our method keeps the model backbone fixed, trains only lightweight routers for head activation, and leverages our custom Select Head Attention kernel to achieve real wall-clock speedups. We believe prior work was limited by the absence of efficient sparse attention primitives, and with the integration of our kernel, these and future approaches have the potential to achieve both high accuracy and meaningful efficiency gains. We hope this unlocks new opportunities and renews momentum in head sparsity research for accelerating LLM inference.
@@ B.1 MLP Activations
Figure 8: OPT and Prosparse Llama Neuron Batch Activations
@@
OPT models, which use ReLU activations, exhibit strong inherent sparsity. As shown in Figure 8, early layers are significantly sparser than deeper ones, a trend consistent across all OPT models. While increasing batch size reduces overall sparsity, the early layers retain enough sparsity to benefit from selective execution. To test the limits of sparsity, we simulated a batch of 10,000 samples and observed that the activations in the initial layers do approach dense and are not, in fact, dead neurons. However, for practical batch sizes, these layers are sufficiently sparse to support accelerated inference. Recent work has shown that sparsity in ReLU-based models tends to increase during pretraining [sparsinglaw]. The OPT models, having undergone extensive large-scale pretraining, likely benefit from this effect, resulting in more structured and persistent activation sparsity.
Figure 8 shows that Prosparse introduces deeper-layer sparsity, but with high variance and similarly degrading trends at larger batch sizes. Unlike OPT, these Sparse-LLaMA variants are only lightly fine-tuned on smaller datasets, which likely limits the emergence of stable sparsity patterns observed in fully pretrained models. Given these observations, we focus our selective execution in LLaMA models on attention head sparsity, where the activation structure remains more consistent across batches.
@@ Appendix C Sparsity Prediction
Our MLP router follows the design used in prior contextual sparsity work [Dejavu, powerinfer]. Each router is a two-layer feedforward network with a hidden dimension of 1024, trained independently for each transformer layer. To partially hide the latency of the MLP router, we overlap its execution with the Attention layer. It is trained using supervised learning with ground-truth neuron activations derived from dense MLP outputs. At inference time, the router produces scalar logits for each neuron, which are used to rank and select the top-$k$ neurons for activation. The attention head routers use a single layer feedforward network as described in Section 4.2. Since the attention routers are much smaller than the MLP routers, we simply run them synchronously before each attention layer. The routers are optimized as binary classifiers using a binary cross-entropy loss, with the AdamW optimizer. We use a batch size of 64, a learning rate of 1e-4, and early stopping over a maximum of 20 epochs. The LLM parameters are frozen during router training. Supervision data is collected from inference runs on the WikiText-2 dataset, as described in Section 5. To determine minimal top-k values for the MLP layers, we apply a simple greedy algorithm (Algorithm 2) that incrementally adjusts the threshold to meet the target recall of 99%. This calibration is performed offline for each OPT model variant.
@@
5:  $\hat{A}\leftarrow R(H)$ $\triangleright$ Predict activations